1. Use the step-by-step method on the circuit below to find the inductor current for all time.

2 ΚΩ t = 0 4 ΚΩ + + 40V Σ12 ΚΩ 20 V 100 mH i,

2. Given the two complex numbers:

2 =-3-j2, 2, =4e/609

a) Express both numbers in rectangular, exponential, and phasor forms;

b) Find the sum, the difference, the product, and the quotient of the numbers.

3. Given the three sinusoids:

x(t)=3cos(@t+45°) y(t)=2cos(ot – 30°) z(t)= 4 sin(at+30°)

a) Find their corresponding phasors using a cosine reference;

b) Does x(t) lead or lag y(t), and by how much?
Answer #1

1 To Calculate Celrrent across inductor 1, we will use fiollowing Reg t L {(t) = [ [0+) - ibo)] e at formula +2.60) 1 w WhereR KR + gov when t>O, Carcuit will look like as shown in dig2 Now thes arcuut goes inlo 4K2 Sleady Dale con dihon at t=00. And3 Pulling call the values inlo en ? 13x103 x + į.lt) (20-5) e - (Bix! *+) - (311091) IDOXID-3 + 5 15 e + 5 Taking 5 common fr10 Z. = 40 160 C 13 from ean Q = 46360° (2 2 3.46 - b= 4 sin Go & + j (3.46) j60 And 4 e So za Rectangular form => Ex ponen h5 from 09 Quotient => V13 33.7 4 260 V13 4 | 33.7-60° = Quotient 0.9 LE 26.3) | Am = 1 3 @ 217) 3 Cos 165 +459) ylt 2 Cos (l